# CharBench: Evaluating the Role of Tokenization in Character-Level Tasks

### v1

@@ Introduction
We present CharBench, a comprehensive suite of evaluation tasks designed to systematically assess the performance of LLMs on character-level reasoning tasks in English. CharBench is designed to test LLMs’ ability to reason over characters, with separate tasks for assessing models’ performance on positional understanding (indexing) and occurrence tracking (counting). We evaluate a diverse set of state-of-the-art language models from both proprietary and open-weight architectures over a range of parameter sizes. CharBench proves to be a substantial challenge, with an overall average accuracy of 50.3% across models. Positional understanding appears to be particularly difficult, with average accuracies of 43.6% and 32.4% in the tasks in this category.
@@
To construct CharBench, [^1] Upon publication, we will share all data and code. we developed a set of modular task templates that represent a variety of character-level reasoning scenarios. These templates can be instantiated with arbitrary words and characters, as illustrated in Figure 1. We focused on tasks that are verifiable and can be mapped directly to simple Python functions, such as count() or index(). To populate the benchmark, we uniformly sampled over 10,000 words from the MiniPile dataset (Kaddour 2023), ensuring balanced coverage across word lengths from 4 to 10 characters. This stratified sampling strategy mitigates potential biases related to word length and frequency.
@@ Results
Find First Occurrence Llama-405B $0.212-$ 0.212 -0.325 $0.277-$ 0.277 0.0040.0040.004 $0.191-$ 0.191 $0.317-$ 0.317
@@
GPT-4o (best-performing) $0.136-$ 0.136 $0.170-$ 0.170 $0.128-$ 0.128 0.0590.0590.059 $0.158-$ 0.158 -0.245
Average $0.185-$ 0.185 -0.306 $0.267-$ 0.267 0.0000.0000.000 $0.160-$ 0.160 $0.230-$ 0.230
@@
Mistral-7B -0.099 0.0090.0090.009 0.0860.0860.086 $0.015-$ 0.015 $0.067-$ 0.067 $0.019-$ 0.019
GPT-4o-mini $0.201-$ 0.201 0.2580.2580.258 0.420 $0.067-$ 0.067 $0.127-$ 0.127 $0.161-$ 0.161
GPT-3.5 $0.135-$ 0.135 0.2950.2950.295 0.468 $0.094-$ 0.094 $0.048-$ 0.048 $0.010-$ 0.010
GPT-4o (best-performing) $0.180-$ 0.180 $0.126-$ 0.126 $0.066-$ 0.066 0.0200.0200.020 $0.185-$ 0.185 -0.259
Average -0.185 0.0710.0710.071 0.1830.1830.183 $0.039-$ 0.039 $0.134-$ 0.134 $0.148-$ 0.148
@@
We compute the intersection-over-union of incorrectly answered questions for each task and pair of models. Figure 5 presents these ratios for the Character Frequency Count and Find Last Occurrence tasks. For the former, overlap between models is negligible, suggesting that errors do not follow a shared pattern. In contrast, the latter exhibits substantial overlap across most models, indicating more consistent failure cases. Notably, GPT-4o is an outlier and shows minimal overlap with the other models.
@@
We find that there is variance across models and tasks. For the find_last_occurrence task, models tend to over-index (which can be useful as a guessing bias), whereas for the count_character_frequency task, models tend to undercount. Bias in other tasks is more moderate and varied. [^4] We present the plots for all tasks in the appendix.
@@
In the evaluation prompt, models are explicitly instructed to treat upper-case and lower-case as different characters for the purpose of the question. We examine whether this aspect of upper/lower case has any effect on model performance. We find that the presence of mixed case characters appears to introduce a measurable bias, yet no a catastrophic one. In fact, for some models and tasks, mixed-case inputs seem to improve performance. [^5] We present these trends in the appendix.

### v2

@@ Introduction
We present CharBench [^1] Code and data - https://github.com/omriuz/CharBench , a comprehensive suite of evaluation tasks designed to systematically assess the performance of LLMs on character-level reasoning tasks in English. CharBench is designed to test LLMs’ ability to reason over characters, with separate tasks for assessing models’ performance on positional understanding (indexing) and occurrence tracking (counting). We evaluate a diverse set of state-of-the-art language models from both proprietary and open-weight architectures over a range of parameter sizes. CharBench proves to be a substantial challenge, with an overall average accuracy of 50.3% across models. Positional understanding appears to be particularly difficult, with average accuracies of 43.6% and 32.4% in the tasks in this category.
@@
To construct CharBench, we developed a set of modular task templates that represent a variety of character-level reasoning scenarios. These templates can be instantiated with arbitrary words and characters, as illustrated in Figure 1. We focused on tasks that are verifiable and can be mapped directly to simple Python functions, such as count() or index(). To populate the benchmark, we uniformly sampled over 10,000 words from the MiniPile dataset (Kaddour 2023), ensuring balanced coverage across word lengths from 4 to 10 characters. This stratified sampling strategy mitigates potential biases related to word length and frequency.
@@ Results
Find First Occurrence Llama-405B 0.212-0.212 -0.325 0.277-0.277 $0.004$ 0.191-0.191 0.317-0.317
@@
GPT-4o (best-performing) 0.136-0.136 0.170-0.170 0.128-0.128 $0.059$ 0.158-0.158 -0.245
Average 0.185-0.185 -0.306 0.267-0.267 $0.000$ 0.160-0.160 0.230-0.230
@@
Mistral-7B -0.099 $0.009$ $0.086$ 0.015-0.015 0.067-0.067 0.019-0.019
GPT-4o-mini 0.201-0.201 $0.258$ 0.420 0.067-0.067 0.127-0.127 0.161-0.161
GPT-3.5 0.135-0.135 $0.295$ 0.468 0.094-0.094 0.048-0.048 0.010-0.010
GPT-4o (best-performing) 0.180-0.180 0.126-0.126 0.066-0.066 $0.020$ 0.185-0.185 -0.259
Average -0.185 $0.071$ $0.183$ 0.039-0.039 0.134-0.134 0.148-0.148
@@
We compute the intersection-over-union of incorrectly answered questions for each task and pair of models. Figure 5 presents these ratios for the Character Frequency Count and Find Last Occurrence tasks. For the former, overlap between models is negligible, suggesting that errors do not follow a shared pattern. In contrast, the latter exhibits substantial overlap across most models, indicating more consistent failure cases. Notably, GPT-4o is an outlier and shows minimal overlap with the other models. Appendix  B shows the overlaps for other tasks.
@@
We find that there is variance across models and tasks. in Appendix D, Figure 11 shows that For the find_last_occurrence task, models tend to over-index (which can be useful as a guessing bias), whereas Figure 8 for the count_character_frequency task, models tend to undercount. Figure 9 and Figure 10 Bias in other tasks is more moderate and varied.
@@
In the evaluation prompt, models are explicitly instructed to treat upper-case and lower-case as different characters for the purpose of the question. We examine whether this aspect of upper/lower case has any effect on model performance. We find that the presence of mixed case characters appears to introduce a measurable bias, yet no a catastrophic one. In fact, for some models and tasks, mixed-case inputs seem to improve performance. Appendix F illustrates these patterns.